The purpose of this question is to compare the variability of x¯¯¯1 and x¯¯¯2 with the variability of (x¯¯¯1−x¯¯¯2). (

a) Suppose the first sample of 100 observations is selected from a population with mean μ1=180 and variance σ21=1180. Construct an interval extending 2 standard deviations of x¯¯¯1 on each side of μ1.

______ ≤ μ 1≤ _______

(b) Suppose the second sample of 100 observations is selected from a population with mean μ2=180 and variance σ22=970. Construct an interval extending 2 standard deviations of x¯¯¯2 on each side of μ

___ ≤ μ2 ≤ ______
Answer #1

a)
lower limit = 180 - 2*sqrt(1180/100) = 173.13
upper limit = 180 + 2*sqrt(1180/100) = 186.87

173.13 < mu1 < 186.87

b)
lower limit = 180 - 2*sqrt(970/100) = 173.77
upper limit = 180 + 2*sqrt(970/100) = 186.23

173.77 < mu2 < 186.23